Leon drove 270 miles to the lodge in the same time as john drove 330 miles to the lodge. If John drove 10 miles per hour faster than Leon, then how fast did each of them drive?
Answer by checkley77(12844)

D=RT
270=RT
T=270/R
330=(R+10)T
330=(R+10)(270/R)
330=(270R+2,700)/R
330R=270R+2,700
60R=2,700
R=2,700/60
R=45 MPH. ANS. FOR THE LEON'S SPEED.
45+10=55 MPH. ANS. FOR JOHN'S SPEED.
PROOF:
330=55(270/45)
330=55*6
330=330